OpenStudy (anonymous):

Solve for x: 16x^2 − 81 = 0 A. x = 4, 9 B. x = −4, −9 C. x = 9 over 4, x = −9 over 4 D. x = 81 over 16, x = −81 over 16

OpenStudy (anonymous):

Factor it you'll get (4x-9)(4x+9)=0 solve for x for both equations, x = 9/4 and x = -9/4